# MAVFlow: Preserving Paralinguistic Elements with Conditional Flow Matching for Zero-Shot AV2AV Multilingual Translation

### v1

@@ Abstract
Despite recent advances in text-to-speech (TTS) models, audio-visual to audio-visual (AV2AV) translation still faces a critical challenge: maintaining speaker consistency between the original and translated vocal and facial features. To address this issue, we propose a conditional flow matching (CFM) zero-shot audio-visual renderer that utilizes strong dual guidance from both audio and visual modalities. By leveraging multi-modal guidance with CFM, our model robustly preserves speaker-specific characteristics and significantly enhances zero-shot AV2AV translation abilities. For the audio modality, we enhance the CFM process by integrating robust speaker embeddings with x-vectors, which serve to bolster speaker consistency. Additionally, we convey emotional nuances to the face rendering module. The guidance provided by both audio and visual cues remains independent of semantic or linguistic content, allowing our renderer to effectively handle zero-shot translation tasks for monolingual speakers in different languages. We empirically demonstrate that the inclusion of high-quality mel-spectrograms conditioned on facial information not only enhances the quality of the synthesized speech but also positively influences facial generation, leading to overall performance improvements.
22footnotetext: Corresponding author
@@ 1 Introduction
At a fundamental level, the transition from single-modality to dual-modality translation is achievable via cascaded approaches. A typical pipeline involves using an automatic speech recognition (ASR) model [3, 51] to transcribe the source audio into text, subsequently applying NMT [18, 15] for language conversion, and finally synthesizing speech via text-to-speech (TTS) systems [7, 63, 8] in conjunction with talking face generation (TFG) models [67, 50, 48]. However, such cascaded methods are complex and often suffer from significant information loss due to repeated modality transformations and intermediate text representations. To overcome these challenges, direct audio-visual-to-audio-visual (AV2AV) translation approaches have been introduced [24, 13]. These methods bypass textual representations by leveraging discrete units obtained from self-supervised multimodal models (e.g., multilingual AV-HuBERT [54, 13]), enabling a more direct and efficient translation of audio-visual source inputs.
Despite these advances, the AV2AV translation still faces a critical challenge: preserving speaker consistency (e.g., tone, pitch, or facial expressions) between the original and translated audio-visual data(as shown in Figure 1). This limitation is exacerbated by the absence of datasets wherein the same speaker articulates identical content in multiple languages, necessitating zero-shot strategies for speaker preservation. Current AV2AV approaches [13, 12] adopt a simple architecture that relies on using speaker embeddings, combining speaker-specific d-vectors [60] within the speech vocoder [29, 50]. This yet underexplored structure of AV2AV, which does not employ advanced conditional generation techniques, still restricts its ability to maintain speaker consistency. Moreover, generating audio and visual components is separated, using a single-modality embedding in each part. This may have limitations from a multimodal perspective since only the reference audio is used, while visual cues could also be considered for speech generation.
In this study, to preserve paralinguistic elements of speakers during the linguistic translation, we propose MAVFlow, a conditional flow matching (CFM, [43]) based zero-shot audio-visual renderer that leverages dual guidance from audio and visual modalities. Notably, for ideal multilingual translation scenarios, a speaker’s voice characteristics and facial information (e.g., appearance or emotion) must remain consistent regardless of language [65, 5, 1]. Based on this hypothesis, we adopt a guidance strategy that utilizes speaker embeddings from audio and emotion embeddings from visual input. This strategy enables the complementary capture of paralinguistic information, which are commonly shared across both audio and visual modalities.
@@
    We empirically demonstrate that our dual guidance improves the consistency of synthesized speech by an average of 36% on the MuAViC dataset [3], while enhancing face generation with gains in lip-sync accuracy (+0.87) and visual quality score (-0.61) on direct system.
@@ 2.1 Spoken Language Translation
Despite advances, S2ST systems primarily focus on audio signals, often neglecting the alignment between translated speech and visual information, which is crucial for cross-lingual scenarios such as video conferencing or dubbing. This can lead to lip-sync mismatches [42, 31], disrupting realistic multimodal experiences. Speech-driven talking face generation (TFG) has been explored to synchronize video with translated speech [32, 62]. More recently, TransFace [12] and AV2AV [13] jointly generate synchronized audio and visual outputs. However, achieving natural cross-lingual experiences remains challenging, particularly in preserving speaker identity and maintaining emotional consistency.
@@ 3.1 Audio-Visual Speech Unit Translation
Recent advances in direct audio-visual to audio-visual translation have leveraged discrete speech units to bypass intermediate text transcription, thereby avoiding delays and error propagation of cascaded systems and expanding applicability [12]. To obtain translated AV units, our system follows two-stage procedure. First, we extract discrete AV units from an input sequence using the unit extractor, m-AVHuBERT [13], which has been pretrained on 7,000 hours of multilingual audio-visual data. Second, we pass the extracted discrete units to a unit-to-unit (U2U) translation module [13], which translates them into the counterparts of target language. The translated units are subsequently converted into intermediate features (mel-spectrograms) via CFM, and finally transformed back into audio-visual form through the vocoder and face decoder. Notably, both the unit extractor and the U2U translation module are identical to those used in our prior work [13], thus ensuring consistency in performance.
3.2 Optimal-Transport CFM
Conditional flow matching (CFM) is a conditional flow-based approach for learning generative models, particularly applied to mel-spectrograms in audio-visual synthesis tasks. Unlike conventional flow-based models, which learn a bijective mapping between a simple prior distribution (e.g., Gaussian noise) and a mel-spectrogram target distribution, CFM directly optimizes the trajectory connecting the two distributions using optimal transport (OT). This facilitates the effective generation of data distributions conditioned on auxiliary information, such as text embedding, audio-visual features, or speaker embeddings, by learning an appropriate conditional vector field [35, 59].
In our framework, data distribution $p(X)$ is connected to a mel-spectrogram representations, which are connected to a noise distribution $\pi(X)$ through a continuous trajectory $\{X_{t}\}_{t=0}^{1}$. Here, $p(X)$ and $\pi(X)$ denote the mel-spectrogram data distribution (target distribution) and the noise prior distribution, respectively. The trajectory optimized by OT-CFM continuously transforms $\pi(X)$ into $p(X)$. In OT-CFM, this trajectory is optimized in the sense of optimal transport. Specifically, for $t\in[0,1]$,
@@ 3.2 Optimal-Transport CFM
and we train a conditional vector field $\nu_{t}$ that integrates both audio and visual features. The OT-CFM optimization objective is defined by
@@
(a) Framework of MAVFlow
(b) Decoder of OT-CFM model
Figure 2: Overall framework and detailed architecture of MAVFlow. (a) An overview of the proposed MAVFlow translation system. (b) U-Net transformer decoder architecture with multimodal guidance $\mathbf{a}_{\mathrm{spk}}$ and $\mathbf{v}_{\mathrm{spk,t}}$ to generated guided mel-spectrogram.
@@ 4 MAVFlow
To effectively preserve speaker-specific characteristics such as voice consistency and facial expressions in multilingual audio-visual translation, we introduce MAVFlow, comprising four main stages: (1) Audio-Visual Speech Unit Translation, which we have outlined in Section 3.1; (2) Duration Length Regulator; (3) Multimodal Guidance; and (4) CFM based AV-Renderer which effectively integrates paralinguistic multimodal guidance with linguistic audio-visual units to synthesize audio-visual outputs. The overall architecture and pipeline of MAVFlow are illustrated in Figure 2.
@@ 4.1 Duration Length Regulator
Since the output of the U2U Translation model is deduplicated, it is necessary to predict and expand the duration of each unit. To achieve this, we employed a Duration Length Regulator, adapting duration prediction concepts previously explored in Text-to-Speech (TTS) synthesis specifically for our audio-visual translation task. The model predicts unit durations using two 1D-convolution layers and a classifier trained with Mean Squared Error (MSE) loss computed in the logarithmic domain against ground truth durations. Sequences are then resampled via linear interpolation to match the batch’s maximum length, with padding masked to retain valid information. During inference, audio-visual units are separately interpolated to their target lengths and combined afterward, maintaining clear modality boundaries and ensuring consistent feature alignment for subsequent synthesis.
@@ 4.2 Multimodal Guidance
$\mathbf{a}_{\mathrm{spk}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{a}_{\mathrm{utt},i}$ (4)
where $\mathbf{a}_{\mathrm{utt},i}$ is the x-vector extracted from the $i$th utterance of a given speaker, and $N$ is the total number of utterances for that speaker. The use of such an averaged speaker embedding allows the model to learn general speaker information during training, thereby enabling the model to robustly learn common speaker traits.
@@
In addition to audio cues, we incorporate paralinguistic speaker face emotional embeddings to ensure the model also learns from visual characteristics of the speaker. We adopt the EmoFAN [58] to extract facial emotion embedding from each frames of the speaker. Specifically, emotional information in a speaker’s utterance can vary dynamically across frames. For instance, a speaker may start smiling partway through an utterance or shift emotional states over time. Thus, distinct emotion embeddings $\mathbf{v}_{\mathrm{spk},t}$ are added as guidance for generating each mel-spectrogram frame $X_{t}$:
$\mathbf{v}_{\mathrm{spk},t}=Emo(\mathbf{f}_{\mathrm{t}})$ (5)
where $\mathbf{v}_{\mathrm{spk},t}$ denotes the face embedding of the $t$-th sampled frame $f_{t}$, and $Emo$ is facial emotion extractor. This reflects a distinctive aspect of the cross-lingual scenario, where frame-level speaker audio characteristics vary according to language-specific phonetic and prosodic differences (e.g., variations in accent, intonation patterns, rhythm, and stress placement), whereas emotional information remains consistent across languages.
4.3 CFM based Zero-Shot AV-Renderer
@@
To effectively synthesize intermediate mel-spectrograms from audio-visual units, MAVFlow incorporates multimodal information to capture paralinguistic features introduce in Section 4.2. Specifically, we employ CFM to guide the mel-spectrogram generation process, optimizing an objective defined as:
@@
where $\phi_{t}^{OT}(X_{0},X_{1})$ is $(1-(1-\sigma)t)X_{0}+tX_{1}$ and $\omega_{t}\left(\phi_{t}^{OT}(X_{0},X_{1})\middle|X_{1}\right)$ is $X_{1}-(1-\sigma)X_{0}$. The multimodal embeddings consist of a global speaker embedding $\mathbf{a}_{\mathrm{spk}}$, uniformly applied across all frames, and a frame-level emotion embedding $\mathbf{v}_{\mathrm{spk,t}}$, dynamically varying per timestep. These embeddings, together with the linguistic speech tokens $\{\mu_{l}\}_{1:L}$ and the masked mel-spectrogram $\tilde{X}_{1}$, are jointly fed into the neural network $N_{\theta}$ to match the conditional vector field parameterized by $\theta$, facilitating the integration of global speaker characteristics and local emotional dynamics (as shown in Figure 2(a)).
$\begin{array}[]{l}\nu_{t}\bigl{(}\phi_{t}^{OT}(X_{0},X_{1})\mid\theta\bigl{)}\\[5.0pt] =N_{\theta}\bigl{(}\phi_{t}^{OT}(X_{0},X_{1}),t;\mathbf{a}_{\mathrm{spk}},\mathbf{v}_{\mathrm{spk,t}},\{\mu_{l}\}_{1:L},\tilde{X}_{1}\bigl{)}\end{array}$ (7)
@@ 5.1 Implementation Details
For training and evaluation, we utilize MuAViC [3], a multilingual audio-visual corpus comprising 1,200 hours of transcribed speech from thousands of speakers, curated from LRS3 [2] and mTEDx [53]. We use five languages: English, Spanish, French, Italian, and Portuguese. Since MuAViC does not contain emotion labels, we employ an additional dataset, CREMA-D [6], for emotion evaluation. CREMA-D consists of 7,442 short video clips featuring 91 adult actors expressing seven different emotions: anger, disgust, fear, happy, neutral, sad, and surprise. Each clip captures an actor uttering a sentence while simultaneously providing facial expressions and vocal information, making it a suitable dataset for evaluating our model’s performance on emotional maintenance.
@@
MAVFlow uses the CFM model pretrained on the LibriTTS [66] as the initial point for more efficient learning. The model operates on 8 RTX A6000 GPUs with a constant learning rate of 0.0001. The speaker embedding and emotional embedding extracted from each audio and visual input—originally 192 and 256 dimensions, respectively—are compressed to an 80-dimensional representation and used as guidance for the OT-CFM. To convert the generated mel-spectrogram into a raw audio waveform, we trained HiFi-GAN [29], model on the LRS3 dataset. We use the same multi-scale L1 and discriminator loss functions proposed in the HiFiGAN. For precise lip-sync and facial expression generation, we use pretrained Wav2Lip [50] on the LRS2 [56] dataset.
@@ 5.2 Baseline Methods
There exist only two systems, AV2AV [13] and Transface [12], that directly utilize units without generating text in the intermediate process. However, our goal is to develop a zero-shot model that generates translated speech while maximally preserving the original speaker’s paralinguistics. Therefore, Transface, which follows a similar approach to AV2AV but does not incorporate additional speaker embeddings—thus not supporting zero-shot audio generation—was excluded from our comparison. Accordingly, for reasonable performance comparison, we establish baselines by combining existing systems in a cascaded manner and compare our proposed method against them. Specifically, the cascaded systems are built based on the latest off-the-shelf pre-trained models such as AVSR [3], ASR [3], AV2T [3], A2T [3], NMT [8], TTS [7, 8], and TFG [50].

### v2

@@ Abstract
Despite recent advances in text-to-speech (TTS) models, audio-visual-to-audio-visual (AV2AV) translation still faces a critical challenge: maintaining speaker consistency between the original and translated vocal and facial features. To address this issue, we propose a conditional flow matching (CFM) zero-shot audio-visual renderer that utilizes strong dual guidance from both audio and visual modalities. By leveraging multimodal guidance with CFM, our model robustly preserves speaker-specific characteristics and enhances zero-shot AV2AV translation abilities. For the audio modality, we enhance the CFM process by integrating robust speaker embeddings with x-vectors, which serve to bolster speaker consistency. Additionally, we convey emotional nuances to the face rendering module. The guidance provided by both audio and visual cues remains independent of semantic or linguistic content, allowing our renderer to effectively handle zero-shot translation tasks for monolingual speakers in different languages. We empirically demonstrate that the inclusion of high-quality mel-spectrograms conditioned on facial information not only enhances the quality of the synthesized speech but also positively influences facial generation, leading to overall performance improvements in LSE and FID score. Our code is available at https://github.com/Peter-SungwooCho/MAVFlow.
@@ 1 Introduction
At a fundamental level, the transition from single-modality to dual-modality translation is achievable via cascaded approaches. A typical pipeline involves using an automatic speech recognition (ASR) model [3, 52] to transcribe the source audio into text, subsequently applying NMT [19, 15] for language conversion, and finally synthesizing speech via text-to-speech (TTS) systems [7, 65, 8] in conjunction with talking face generation (TFG) models [69, 51, 49]. However, such cascaded methods are complex and often suffer from significant information loss due to repeated modality transformations and intermediate text representations. To overcome these challenges, direct audio-visual-to-audio-visual (AV2AV) translation approaches have been introduced [25, 13, 12]. These methods bypass textual representations by leveraging discrete units obtained from self-supervised multimodal models (e.g., multilingual AV-HuBERT [55, 13]), enabling a more direct and efficient translation of audio-visual source inputs.
Despite these advances, the AV2AV translation still faces a critical challenge: preserving speaker consistency (e.g., tone, pitch, or facial expressions) between the original and translated audio-visual data (as shown in Figure 1). This limitation is exacerbated by the absence of datasets wherein the same speaker articulates identical content in multiple languages, necessitating zero-shot strategies for speaker preservation. Current AV2AV approaches [13, 12] adopt a simple architecture that relies on using speaker embeddings, combining speaker-specific d-vectors [62] within the speech vocoder [30, 51]. This yet underexplored structure of AV2AV, which does not employ advanced conditional generation techniques, still restricts its ability to maintain speaker consistency. Moreover, generating audio and visual components are separated, using a single-modality embedding in each part. This has limitations from a multimodal perspective since only the reference audio is used, while visual cues could also be considered for speech generation.
In this study, to preserve paralinguistic elements of speakers during the linguistic translation, we propose MAVFlow, a Conditional Flow Matching (CFM) [44] based zero-shot audio-visual renderer that leverages dual guidance from audio and visual modalities. Notably, for ideal multilingual translation scenarios, a speaker’s voice characteristics and facial information (e.g., appearance or emotion) must remain consistent regardless of language [67, 5, 1]. Based on this hypothesis, we adopt a guidance strategy that utilizes speaker embeddings from audio and emotion embeddings from visual input. This strategy enables the complementary capture of paralinguistic information, which are commonly shared across both audio and visual modalities.
@@
    We empirically demonstrate that our dual guidance improves the consistency of speaker identity in synthesized speech by an average of 36% on the MuAViC dataset [3], while enhancing face generation with gains in lip-sync accuracy ($+$0.87) and visual quality score (--0.61) on textless system.
@@ 2.1 Spoken Language Translation
Despite advances, S2ST systems primarily focus on audio signals, often neglecting the alignment between translated speech and visual information, which is crucial for cross-lingual scenarios such as video conferencing or dubbing. This can lead to lip-sync mismatches [43, 32], disrupting realistic multimodal experiences. Speech-driven TFG has been explored to synchronize video with translated speech [33, 64]. More recently, TransFace [12] and AV2AV [13] jointly generate synchronized audio and visual outputs. However, achieving natural cross-lingual experiences remains challenging, particularly in preserving speaker identity and maintaining emotional consistency.
@@ 3.1 Audio-Visual Speech Unit Translation
Recent advances in direct audio-visual-to-audio-visual translation have leveraged discrete speech units to bypass intermediate text transcription, thereby avoiding delays and error propagation of cascaded systems and expanding applicability [12]. To obtain translated AV units, our system follows two-stage procedure. First, we extract discrete AV units from an input sequence using the unit extractor, m-AVHuBERT [13], which has been pretrained on 7,000 hours of multilingual audio-visual data. Second, we pass the extracted discrete units to a unit-to-unit (U2U) translation module [13], which translates them into the counterparts of target language. The translated units are subsequently converted into intermediate features (mel-spectrograms) via CFM, and finally transformed back into audio-visual form through the vocoder and face decoder. Notably, both the unit extractor and the U2U translation module are identical to those used in our prior work [13], thus ensuring consistency in performance.
3.2 Optimal Transport CFM
Conditional Flow Matching (CFM) is a framework that leverages conditional flows to train generative models, particularly applied to generate mel-spectrograms in audio synthesis tasks [44, 17]. Unlike conventional flow-based models, which learn a bijective mapping between a simple prior distribution (e.g., Gaussian noise) and a mel-spectrogram target distribution, CFM directly optimizes the trajectory connecting the two distributions using optimal transport (OT). This facilitates the effective generation of data distributions conditioned on auxiliary information, such as text embedding, audio-visual features, or speaker embeddings, by learning an appropriate conditional vector field [36, 61].
In our framework, data distribution $p(X)$ is connected to a mel-spectrogram representations, which are connected to a noise distribution $\pi(X)$ through a continuous trajectory $\{X_{t}\}_{t=0}^{1}$. Here, $p(X)$ and $\pi(X)$ denote the mel-spectrogram data distribution (target distribution) and the noise prior distribution, respectively. The trajectory that continuously transforms $\pi(X)$ into $p(X)$, which is optimized by OT-CFM in the sense of optimal transport. Specifically, for $t\in[0,1]$,
@@ 3.2 Optimal Transport CFM
and we train a conditional vector field $\nu_{t}$ that integrates both audio and visual features. The OT-CFM optimization objective function is defined by
@@
(a) Overall framework of MAVFlow
(b) OT-CFM decoder structure
Figure 2: Overall framework and detailed architecture of MAVFlow. (a) An overview of the proposed MAVFlow translation system. (b) OT-CFM’s Transformer decoder structure with multimodal guidance $\mathbf{a}_{\mathrm{spk}}$ and $\mathbf{v}_{\mathrm{spk,t}}$ to generate guided mel-spectrogram.
@@ 4 MAVFlow
To effectively preserve speaker-specific characteristics such as voice consistency and facial expressions in multilingual audio-visual translation, we introduce MAVFlow, comprising four main stages: (i) Audio-Visual Speech Unit Translation, which we have outlined in Section 3.1; (ii) Duration Length Regulator; (iii) Multimodal Guidance; and (iv) CFM-based Zero-Shot AV-Renderer which effectively integrates paralinguistic multimodal guidance with linguistic audio-visual units to synthesize audio-visual outputs. The overall architecture and pipeline of MAVFlow are illustrated in Figure 2.
@@ 4.1 Duration Length Regulator
Since the output of the U2U translation module is deduplicated, it is necessary to predict and expand the duration of each unit. To achieve this, we employ a Duration Length Regulator, adapting duration prediction concepts previously explored in TTS synthesis specifically for our audio-visual translation task. We adopt a similar duration prediction structure and loss function from AV2AV [13], using two 1D-convolution layers with a classifier, where the objective function is the MSE loss in the log domain. However, our Duration Length Regulator differs in that it interpolates the generated audio to match the length of the original source audio. This design addresses a critical constraint in real-world movie dubbing scenarios, where the video length must remain consistent before and after translation—an aspect not considered in AV2AV.
@@ 4.2 Multimodal Guidance
$\mathbf{a}_{\mathrm{spk}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{a}_{\mathrm{utt},i}\vskip-1.0pt$ (4)
where $\mathbf{a}_{\mathrm{utt},i}$ is the x-vector extracted from the $i$-th utterance of a given speaker, and $N$ is the total number of utterances for that speaker. The use of such an averaged speaker embedding allows the model to learn general speaker information during training, thereby enabling the model to robustly learn common speaker traits.
@@
In addition to audio cues, we incorporate paralinguistic speaker face emotional embeddings to ensure that the model also learns from visual characteristics of the speaker. We adopt EmoFAN [60] to extract facial emotion embeddings from each frame. Specifically, emotional information in a speaker’s utterance can vary dynamically across frames. For instance, a speaker may start smiling partway through an utterance or shift emotional states over time. Thus, distinct emotion embeddings $\mathbf{v}_{\mathrm{spk},t}$ are added as guidance for generating each mel-spectrogram frame $X_{t}$:
$\mathbf{v}_{\mathrm{spk},t}=\text{Emo}(\mathbf{f}_{\mathrm{t}})$ (5)
[…]
4.3 CFM-based Zero-Shot AV-Renderer
@@
To effectively synthesize intermediate mel-spectrograms from audio-visual units, MAVFlow incorporates multimodal information to capture paralinguistic features, as introduced in Section 4.2. Specifically, we employ CFM to guide the mel-spectrogram generation process, optimizing an objective defined as:
@@
where $\phi_{t}^{OT}(X_{0},X_{1})$ is $(1-(1-\sigma)t)X_{0}+tX_{1}$ and $\omega_{t}\left(\phi_{t}^{OT}(X_{0},X_{1})\middle|X_{1}\right)$ is $X_{1}-(1-\sigma)X_{0}$.
The multimodal embeddings consist of a global speaker embedding $\mathbf{a}_{\mathrm{spk}}$, uniformly applied across all frames, and a frame-level emotion embedding $\mathbf{v}_{\mathrm{spk,t}}$, dynamically varying per timestep. These embeddings, together with the linguistic speech tokens $\{\mu_{l}\}_{1:L}$ and the masked mel-spectrogram $\tilde{X}_{1}$, are jointly fed into the neural network $N_{\theta}$ to match the conditional vector field parameterized by $\theta$, facilitating the integration of global speaker characteristics and local emotional dynamics (as shown in Figure 2(a)).
$\begin{array}[]{l}\nu_{t}\bigl{(}\phi_{t}^{OT}(X_{0},X_{1})\mid\theta\bigl{)}\\[5.0pt] \hskip-8.5359pt=N_{\theta}\bigl{(}\phi_{t}^{OT}(X_{0},X_{1}),t;\mathbf{a}_{\mathrm{spk}},\mathbf{v}_{\mathrm{spk,t}},\{\mu_{l}\}_{1:L},\tilde{X}_{1}\bigl{)}\end{array}$ (7)
@@ 5.1 Implementation Details
For training and evaluation, we utilize MuAViC [3], a multilingual audio-visual corpus comprising 1,200 hours of transcribed speech from thousands of speakers, curated from LRS3 [2] and mTEDx [54]. We use five languages: English, Spanish, French, Italian, and Portuguese. Since MuAViC does not contain emotion labels, we employ an additional dataset, CREMA-D [6], for emotion evaluation. CREMA-D consists of 7,442 short video clips featuring 91 adult actors expressing six different emotions: anger, disgust, fear, happy, neutral, and sad. Each clip captures an actor uttering a sentence while simultaneously providing facial expressions and vocal information, making it a suitable dataset for evaluating our model’s performance on emotional maintenance.
@@
MAVFlow uses the CFM model pretrained on the LibriTTS [68] as the initial point for more efficient learning. The model is trained on 8 RTX A6000 GPUs with a constant learning rate of 0.0001. The speaker embedding and emotional embedding extracted from each audio and visual input—originally 192 and 256 dimensions, respectively—are compressed to an 80-dimensional representation and used as guidance for the OT-CFM. To convert the generated mel-spectrogram into a raw audio waveform, we train HiFi-GAN [30] on the LRS3 dataset. We use the same multi-scale L1 and discriminator loss functions proposed in HiFi-GAN. For precise lip-sync and facial expression generation, we use pretrained Wav2Lip [51] on the LRS2 [57] dataset. Details about the inference time are provided in Appendix C.
@@ 5.2 Baseline Methods
There exist only two textless systems, AV2AV [13] and Transface [12], that directly utilize units without generating text in the intermediate process. However, our goal is to develop a zero-shot model that generates translated speech while maximally preserving the original speaker’s paralinguistics. Therefore, Transface, which follows a similar approach to AV2AV but does not incorporate additional speaker embeddings—thus not supporting zero-shot audio generation—was excluded from our comparison. Accordingly, for reasonable performance comparison, we establish baselines by combining existing systems in a cascaded manner and compare our proposed method against them. Specifically, the cascaded systems are built based on the latest off-the-shelf pre-trained models such as AVSR [3], ASR [3], AV2T [3], A2T [3], NMT [8], TTS [7, 8], and TFG [51].